OpenStudy (anonymous):

Find the standard form of the equation of the parabola with the given characteristic and vertex at the origin. Focus: (1/2,0)

OpenStudy (mertsj):

Opens to the right. p = 1/2 y^2=4px

OpenStudy (anonymous):

so the equation would be (x-h)^2=4p(y-k) right?